OpenStudy (anonymous):

A is 2x2 matrix and its eigenvalue mu with eigenvector v. P(A)= x^2+ 5x. How do we show that v is the eigenvector and what is its related eigenvalue?

OpenStudy (e.mccormick):

\[A\vec{v}=\mu \vec{v}\]

OpenStudy (anonymous):

do we not do anything with P(A)? like with it being equal to x^2+5x?

OpenStudy (e.mccormick):

Well, need to get something useable out of it. Just pointing out one of the ways to prove an eigenvalue/eigenvector.

OpenStudy (anonymous):

I do not think this is very explanatory..

OpenStudy (e.mccormick):

Well, what does \(P(A)= x^2+ 5x.\) tell you about A?

OpenStudy (anonymous):

Thats what I do not get... I do not know how it is connected to the question and how to use the equation... bec I know the general formula to descrive P(A)..

OpenStudy (anonymous):

P(A)= A^2+5A..

OpenStudy (e.mccormick):

Yah, I'm not seeing a reference to that in what I have. Sorry I do not think I can help clarify it. Which I could!

OpenStudy (anonymous):

I am making a correction-- P(A)= A^2+5A not to x^2+5x

OpenStudy (anonymous):

@KingGeorge

OpenStudy (kinggeorge):

Well, if P(x)=x^2+5x, then the roots of this are the eigenvalues. Since \(x^2+5x=x(x+5)\), the eigenvalues are \(0,-5\). At least that's what it is if I remember my definitions of P(A) correctly.

OpenStudy (kinggeorge):

Do you have actual numbers for A and v?

OpenStudy (e.mccormick):

That is what I suspected, opened my book.... and found no reference. Looked online, found no reference... but I have learned that Linear is not as rigorously standard as say Algebra 1, so there is more room for different notations, etc.

OpenStudy (anonymous):

no we do not.. it is just asking How do we show that v is the eigenvector...

OpenStudy (anonymous):

how you found eigenvalue made so much sense.. but just is there anyway that I am supposed to connect it to showing that v is the eigenvector?

OpenStudy (kinggeorge):

Without knowing any values for v or A, I don't know how to continue.

OpenStudy (e.mccormick):

\[A\vec{v}=\mu \vec{v}\implies A\vec{v}=0 \vec{v} \text{ or }A\vec{v}=-5 \vec{v}\]but it relies on having other values.... as KG points out.

OpenStudy (e.mccormick):

Found a reference to the characteristic polynomial that I had not seen before. It is related to the Cayley–Hamilton theorem. \[p(A)=A^2-(a+d)A+(ad-bc)I_2=\left[ \begin{matrix} 0&0\\0&0 \end{matrix} \right]\]

OpenStudy (e.mccormick):

With \(P(A)= x^2+ 5x\) the determinant would be 0 if I match it up with that form. And \(a+d=-5\). Don't know if that gets it any furthur.

OpenStudy (kinggeorge):

That's a good observation. But a+d=5, not -5.

OpenStudy (e.mccormick):

Note the negative in front.

OpenStudy (kinggeorge):

Oh. Right. I accounted for one too many negatives.

OpenStudy (e.mccormick):

\[A^2= \left[ \begin{matrix} a&b\\c&d \end{matrix} \right] \left[ \begin{matrix} a&b\\c&d \end{matrix} \right]\\ \implies A^2= \left[ \begin{matrix} a^2+bc&ab+bd\\ca+dc&cb+d^2\end{matrix} \right]\\ \implies A^2= \left[ \begin{matrix} a^2+bc&-5b\\-5c&cb+d^2\end{matrix} \right]\]

OpenStudy (e.mccormick):

And the determinant of that should still be 0.

OpenStudy (kinggeorge):

Then if we add 5A, and use the Cayley-Hamilton theorem, we also find that \(a^2+5a+bc=d^2+5d+bc=0\).

OpenStudy (e.mccormick):

\((a^2+bc)(cb+d^2)-25=0\)

OpenStudy (e.mccormick):

So two equations that both =0. Too bad Melisac went offline. I think we finally got somewhere. Not far, but something.

OpenStudy (kinggeorge):

If we assume A is diagonal we could get a little farther, but I'm not sure how that would change things.

OpenStudy (e.mccormick):

det[A]=0... does that preclude diagonalizable?

OpenStudy (e.mccormick):

It has to be at least linearly dependent or a row/col of 0 or something like that.

OpenStudy (kinggeorge):

Oops. I meant upper/lower-triangular. But if that's the case, then at least one row or one column must be 0.

OpenStudy (kinggeorge):

But in any case, since A has distinct eigenvalues, it should be diagonalizable to \[\begin{bmatrix}0&0\\0&-5\end{bmatrix}.\](This would be the Jordan canonical form).

OpenStudy (e.mccormick):

AH HA! Yah, you came to what I was just thinking! There has to be a -5 in either a or d, it can not have a value in both!

OpenStudy (kinggeorge):

I don't know if that's necessarily true yet, but it's certainly similar to such a matrix, so we might be able to work with that.

OpenStudy (e.mccormick):

We know it is 2x2 and it has ... ah, yah... they could be a multiple... so one possibility, if they are 0 in one row or colume would mean -5 in only one, but the other possibility, say row 2 is 1/2 row a, leaves that open.

OpenStudy (kinggeorge):

For example \[\begin{bmatrix}1&3\\-2&-6\end{bmatrix}\]would be such a matrix. However, the column vectors are not linearly independent.

OpenStudy (e.mccormick):

yah.\[\left[ \begin{matrix} -10&10\\-5&5 \end{matrix} \right]\]Looks possible too.So lots of options.

OpenStudy (kinggeorge):

Well, this has been interesting, but I've got to sleep now.

OpenStudy (e.mccormick):

yah, we had a good stab at it. I hope Melisac finds something useful in here.